OpenStudy (anonymous):

A rope pulls on a metal box at an angle of 40 degrees with a force of 255 N. The box moves horizontally for 15.0 m. How much work was done on the box? A. 255 J B. 2,930 J C. 2,460 J D. 3,830 J

OpenStudy (michele_laino):

as in a previous exercise, we have to apply the subsequent formula: \[W = Fd\cos \left( {40} \right) = 255 \times 15 \times \cos 40\]

OpenStudy (anonymous):

2930 J is our solution?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! thank you:)

OpenStudy (michele_laino):

:)